# Representing Sugihara monoids

### v3

@@ 1 Introduction
We will view Sugihara monoids as examples of distributive involutive FL-algebras (DInFL-algebras). See Section 2.1 for a precise definition of these algebras. DInFL-algebras have two order-reversing operations $\sim$ and --. If, for all $a\in\mathbf{A}$, ${\sim}a=-a$ then the algebra $\mathbf{A}$ is said to be cyclic. The use of a non-identity order isomorphism $\alpha$ allows for the construction of concrete non-cyclic DInFL-algebras (see item (ii) of Theorem 2.7). Our representation of Sugihara monoids is part of a more general method which can also handle non-cyclic DInFL-algebras.
[…]
@@ 2.1 Sugihara monoids and DInFL-algebras
A residuated lattice is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,\backslash,/\rangle$ such that $\langle A,\wedge,\vee\rangle$ is a lattice, $\langle A,\cdot,1\rangle$ is a monoid and the following holds for all $a,b,c\in A$:
$a\cdot b\leqslant c\quad\Longleftrightarrow\quad a\leqslant c/b\quad\Longleftrightarrow\quad b\leqslant a\backslash c.$
[…]
@@
$a_{i}\to a_{j}=\begin{cases}{\sim}a_{i}\vee a_{j}&\text{if }i\leqslant j\\ {\sim}a_{i}\wedge a_{j}&\text{if }i>j.\end{cases}$
@@
[…]
@@ 2.3 Constructing InFL-algebras of relations
[…]
@@
[…]
@@
The next lemma shows that if $\alpha$ is an order automorphism of a poset $\langle X,\leqslant\rangle$ with an equivalence relation $E$ such that ${\leqslant}\subseteq E$, then $\alpha$ can be used to define an element of $\mathsf{Up}(\mathbf{E})$ to act as a 00 that will give rise to an InFL-algebra.
@@
We will now add $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ to the signature of the distributive residuated lattice $\left\langle\mathsf{Up}\left(\mathbf{E}\right),\cap,\cup,\circ,\leqslant,\backslash,/\right\rangle$ to obtain the FL-algebra $\left\langle\mathsf{Up}\left(\mathbf{E}\right),\cap,\cup,\circ,\leqslant,0,\backslash,/\right\rangle$. The next lemma gives an alternative way of calculating ${\sim}R$ and $-R$ if we set $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$.
@@
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset with $E$ an equivalence relation and $\alpha:X\rightarrow X$ an order automorphism of $\mathbf{X}$ where ${\leqslant}\subseteq E$ and $\alpha\subseteq E$. If $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ then ${\sim}R=R^{c\smile}\circ\alpha$ and $-R=\alpha\circ R^{c\smile}$ for all $R\in\mathsf{Up}\left(\langle X^{2},\preccurlyeq\rangle\right)$.
[…]
@@
[…]
@@
[…]
@@
These relations are upsets of $\langle E,\preccurlyeq\rangle$ and form a Sugihara chain of length $22n+$ 2, i.e.,
@@ 3 Constructing finite odd Sugihara chains
[…]
@@
To prove the left-to-right inclusion, assume $x^{b}\leqslant_{X}y^{d}$. Then $x^{b}=y^{d}$ or $x<y$. If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{d}\right)\in 0$. On the other hand, if $x<y$, then $y\not\leqslant x$, so $y^{-d}\not\leqslant x^{b}$. Again we have $(x^{b},y^{d})\in 0$. For the other inclusion, let $\left(x^{b},y^{d}\right)\in 0=\left(\leqslant_{X}\right)^{c\smile}\circ\alpha$. Then we have $y^{-d}\not\leqslant x^{b}$. If $b=-d$, then $y\not\leqslant x$, so $x<y$. Hence, $x^{b}\leqslant_{X}y^{d}$. On the other hand, if $b\neq-d$, then $y^{-d}\not\leqslant x^{b}$ implies $y\not<x$. This means $x\leqslant y$, and so, since $d=b$, we get $x^{b}\leqslant_{X}y^{d}$.
@@
[…]
@@
[…]
@@
Both $X^{2}$ and $\varnothing$ are upsets of $\left\langle X^{2},\preccurlyeq\right\rangle$. The fact that $\leqslant_{X}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ follows from Lemma 2.4. It would also follow from Lemmas 2.3 and 2.6 that $R_{i}$ is an upset for $i\in\left\{1,\ldots,n\right\}$, if we can show that $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ when $-n\leqslant i<0$.
@@
Notice that when defining $R_{i}$ for $i\in\left\{1,\ldots,n+1\right\}$ we chose to use the unary order-reversing operation $\sim$ instead of --. Since $\alpha\neq\text{id}_{X}$ we know from Theorem 2.7 that the algebra will not be cyclic and hence ${\sim}R\neq-R$ in general. However, the following lemma shows that for the relations $R_{i}$, we will in fact get ${\sim}R_{i}=-R_{i}$.
@@
For each $i\in\left\{-n-1,\ldots,n+1\right\}$, we have $-R_{i}={\sim}R_{i}$.
@@
Case 1: $-n-1\leqslant i<0$. Let $\left(p^{b},q^{d}\right)\notin-R_{i}$. By Lemma 2.2, $\left(p^{b},q^{d}\right)\in\left(-R_{i}\right)^{c}=\left(\alpha\circ\left(R_{i}\right)^{c\smile}\right)^{c}=\alpha\;\circ\;R^{\smile}_{i}$. Therefore, $\left(q^{d},p^{-b}\right)\in R_{i}$, and so $\left(q^{d},p^{-b}\right)\in U_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. This implies $(q,p)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. It thus follows that $\left(q^{-d},p^{b}\right)\in U_{k}\subseteq R_{i}$, i.e., $\left(\alpha\left(q^{d}\right),p^{b}\right)\in U_{k}\subseteq R_{i}$. Hence, $\left(p^{b},\alpha\left(q^{d}\right)\right)\in R_{i}^{\smile}$, and therefore $\left(p^{b},\alpha\left(q^{d}\right)\right)\in\left(\left(R_{i}\right)^{c\smile}\right)^{c}$. From this we get $\left(p^{b},q^{d}\right)\in\left(\left(R_{i}\right)^{c\smile}\right)^{c}\circ\alpha=\left(\left(R_{i}\right)^{c\smile}\circ\alpha\right)^{c}=\left({\sim}R_{i}\right)^{c}$, i.e., $\left(p^{b},q^{d}\right)\notin{\sim}R_{i}$.
@@
Case 2: $10<i\leqslant n+$ 1. This case follows from the previous case and the fact that -- and $\sim$ satisfy the involutive property.
@@
This is clearly true for $i=0$ and $i=n+1$. Since $\varnothing$ is vacuously transitive, the statement also holds for $i=-n-1$. We now consider the following two cases:
Case 1: $-n\leqslant i<0$. Let $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. Then there exist $j,k\in\left\{1,\ldots,n+1+i\right\}$ such that $(p,r)\in L_{j}$ and $(r,q)\in L_{k}$. Hence, by Lemma 3.3, we have $\left(p,q\right)\in L_{\text{min}\{j,k\}}$, and so $\left(p^{b},q^{d}\right)\in U_{\text{min}\{j,k\}}\subseteq R_{i}$.
@@
If $i=-n-1$ (i.e. $R_{i}=\varnothing$) then we clearly have $R_{i}\circ R_{j}=R_{i}$. The case for $j=0$ follows from Lemma 2.2. So we consider the following cases:
Case 1: $-n-1<i<j<0$. Since $R_{j}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{i}$. Let $\left(p^{b},q^{d}\right)\in R_{i}$. Hence, there exists $k\in\left\{1,\ldots,n+1+i\right\}$ such that $\left(p^{b},q^{d}\right)\in U_{k}$, i.e., there exists $k\in\left\{1,\ldots,n+1+i\right\}$ such that $\left(p,q\right)\in L_{k}$. There are two cases:
@@
Case 2: $-n-1<i<0$, $j>0$. Since ${\leqslant_{X}}\subseteq R_{j}$, we get $R_{i}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. From the first part we get $\left(p^{b},r^{e}\right)\in U_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$, which means $(p,r)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. The second part gives $\left(r^{e},q^{d}\right)\in{\sim}R_{-j}=\left(R_{-j}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},r^{b}\right)\notin R_{-j}$. This implies that $\left(q^{-d},r^{e}\right)\notin U_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-j\right\}$. That is, $q_{1}\not<r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $q_{\ell}\not<r_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we get $r_{1}\leqslant q_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $r_{\ell}\leqslant q_{1}$. If $p_{1}<r_{1}$, then $p_{1}<r_{1}\leqslant q_{1}$, and so $\left(p,q\right)\in L_{1}$, i.e., $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}$. If $r_{1}<q_{1}$, we have $p_{1}\leqslant r_{1}<q_{1}$, so again $\left(p,q\right)\in L_{1}$, i.e., $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}$. If $p_{1}=r_{1}=q_{1}$, then it must be the case that $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}$.
Case 3: $i>0$, $-n-1<j<0$. Since $R_{j}\subseteq{\leqslant_{X}}$ we get $R_{i}\circ R_{j}\subseteq R_{i}$. Let $\left(p^{b},q^{d}\right)\in R_{i}$. Then $\left(p^{b},q^{d}\right)\in{\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},p^{b}\right)\notin R_{-i}$. This means that $\left(q^{-d},p^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. That is, $q\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{k-1}\right)=\left(q_{1},\ldots,q_{k-1}\right)$ then $q_{k}\not<p_{k}$. We therefore obtain $p_{1}\leqslant q_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{k-1}\right)=\left(q_{1},\ldots,q_{k-1}\right)$ then $p_{k}\leqslant q_{k}$. Suppose $p_{1}<q_{1}$. By the density of $\mathbb{Q}$, there is some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Now let $r=\left(p_{1},0,\ldots,0\right)$. Then $(p,r)\in L_{1}$ and $(r,q)\in L_{1}$, i.e., $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{-i}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{-i}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. On the other hand, suppose $p_{1}=q_{1}$. We have two cases:

### v4

@@ 1 Introduction
We will view Sugihara monoids as examples of distributive involutive FL-algebras (DInFL-algebras). See Section 2.1 for a precise definition of these algebras. DInFL-algebras have two order-reversing operations $\sim$ and -. If, for all $\in\mathbf{A}, a=a{\sim}$a=-a then the algebra $\mathbf{A}$ is said to be cyclic. The use of a non-identity order isomorphism $\alpha$ allows for the construction of concrete non-cyclic DInFL-algebras (see item (ii) of Theorem 2.7). Our representation of Sugihara monoids is part of a more general method which can also handle non-cyclic DInFL-algebras.
Before giving an outline of the content of this paper, we comment on the recent representation of Sugihara monoids given by Kramer and Maddux [14]. That paper provides an alternative solution to the problem of defining an involutive order-reversing operation on relations that has a fixed point. The Sugihara chains are embedded into proper relations algebras (see [14, Definition 4] for the definition of the binary relations that will be the elements of the Sugihara chain). For an abstract relation algebra, they define relativized subreducts [14, Definition 8] where certain operations from the original relation algebra are retained, while others are replaced by new term-definable operations. This notion of relativized subreduct is then implemented on algebras of relations (see [14, Theorem 2]) in order to obtain a new $\sim$} operation that will have a fixed point, hence allowing for the representation of odd Sugihara monoids. A difference between our approach and the one in [14, Theorem 5] is that our representations are embedded into distributive InFL-algebra of upsets (see Section 2). In their case, the concrete odd Sugihara chain is embedded into a proper relation algebra. A feature of our construction is that the relations which are the elements of the representation are all weakening relations with respect to a partial order.
@@ 2.1 Sugihara monoids and DInFL-algebras
A residuated lattice is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,\backslash$$\rangle$ such that $\langle A,\wedge,\vee\rangle$ is a lattice, $\langle A,\cdot,1\rangle$ is a monoid and the following holds for all $a,b,c\in A$:
$\c.a\cdot b\leqslant c\quad\Longleftrightarrow\quad a\leqslant c/b\quad\Longleftrightarrow\quad b\leqslant a\backslash$ c.
If the monoid operation of a residuated lattice $\mathbf{A}$ is commutative, then both residuals $\backslash$ and // (sometimes denoted $\to$ and $\leftarrow$) coincide. If $a\cdot a=a$ is satisfied for all $a\in A$, then $\mathbf{A}$ is an idempotent residuated lattice. A Sugihara monoid is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot\to,1,{\sim}\rangle$ such that $\langle A,\wedge,\vee,\cdot\to,1\rangle$ is a commutative distributive idempotent residuated lattice, and for all $a,b\in A$ the identities ${\sim}{\sim}a=a$ and $a\to{\sim}b=b\to{\sim}a$ are satisfied. (A commutative residuated lattice satisfying those two identities is often called an involutive commutative residuated lattice.)
It was shown by Dunn [1] that the variety of Sugihara monoids is generated by algebras that are linearly ordered, and thus the linear ordered algebras are of particular interest. For our explicit representations in Section 3, we will only be considering finite Sugihara chains. These are algebras $\mathbf{S}_{n}$ for $n\in\omega$ and $n\geqslant 2$. If $n=2k$ for $k>0$ then $S_{n}=\{a_{-k},\dots,a_{-1},a_{1},\dots,a_{k}\}$, and if $n=2k+1$ for $k>0$ then $S_{n}=\{a_{-k},\dots,a_{-1},a_{0},a_{1},\dots a_{k}\}$. The meet and join are defined as expected: $a_{i}\wedge a_{j}=a_{\text{min}\{i,j\}}$ and $a_{i}\vee a_{j}=a_{\text{max}\{i,j\}}$. The unary operation $\sim$} is defined by ${\sim}a_{j}=a_{-j}$. The monoid operation can be defined as:
@@
[…]
@@
The concrete algebras we describe in Section 2.3 will be special residuated lattices, which we describe briefly here. A Full Lambek algbebra (cf. [11]) is a residuated lattice $\mathbf{L}$ that has an additional constant 0 (about which nothing is assumed) in the signature. We use the common abbreviation FL-algebra. Using the residuals, two unary operations $\sim$} and - can be defined. These are known as linear negations and are defined for $a\in L$ by ${\sim}a=a\backslash 0$ and a=0/a-a=0/a. If $\sim$} and - satisfy the involutive property, i.e. ${\sim}a=a={\sim}{-}a$ for all $a\in L$, then $\langle L,\wedge,\vee,\cdot,1,0,\backslash$$\rangle$ is an InFL-algebra. When the underlying lattice is distributive, we use the abbreviation DInFL-algebra. If ${\sim}a=-a$ for all $a\in L$ then it is a cyclic InFL-algebra.
We prefer to use the linear negations rather than the residuals in the signature as it is known that such an equivalent formulation is available [8, Lemma 2.2]. In addition, that equivalent formulation allows us to leave the 0 out of the signature. That is, we will consider DInFL-algebras as algebras of the following signature: $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,{\sim},$-$\rangle$. Notice that if a Sugihara monoid uses an alternative signature with $\sim$} then it will be a DInFL-alegbra. Moreover, it will be a commutative cyclic idempotent DInFL-algebra.
@@ 2.3 Constructing InFL-algebras of relations
In previous work [3, Section 3], we used the construction outlined below to produce concrete algebras of binary relations which formed distributive quasi relation algebras (DqRAs). These algebras were described by Galatos and Jipsen [8] and are algebras of the form $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,0,{\sim},-,^{\prime}\rangle$. The unary operations $\sim$},- and are dual lattice isomorphisms satisfying some additional conditions. We do not go into further detail here about the definition of quasi relation algebras as we will not make use of the full signature of these algebras. The construction outlined below will produce distributive involutive FL-algebras.
@@
We note that when $E=X^{2}$ the lattice constructed above is identical to that used by Galatos and Jipsen [9, 10], and more recently by Jipsen and Šemrl [13]. In order to represent odd Sugihara monoids as algebras of relations, we will make use of an order automorphism $\alpha$ on the poset $\mathbf{X}$. It is this $\alpha$ that will allow us to define a unary operation $\sim$} on $\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ and an identity $1\in\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ such that ${\sim}1=1$.
@@
The next lemma shows that if $\alpha$ is an order automorphism of a poset $\langle X,\leqslant\rangle$ with an equivalence relation $E$ such that ${\leqslant}\subseteq E$, then $\alpha$ can be used to define an element of $\mathsf{Up}(\mathbf{E})$ to act as a 0 that will give rise to an InFL-algebra.
@@
[…]
@@
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset with $E$ an equivalence relation and $\alpha:X\rightarrow X$ an order automorphism of $\mathbf{X}$ where ${\leqslant}\subseteq E$ and $\alpha\subseteq E$. If $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ then ${\sim}R=R^{c\smile}\circ\alpha$ and $R=\alpha\circ R^{c\smile}$ for all $R\in\mathsf{Up}\left(\langle X^{2},\preccurlyeq\rangle\right)$.
Using the alternative method of calculating $\sim$} and {-} from Lemma 2.6, greatly simplifies the proof of the main theorem of this background section.
@@
Consider the three-element Sugihara chain $\mathbf{S}_{3}$. The underlying set is given by $S_{3}=\{a_{-1},a_{0},a_{1}\}$ where $a_{0}$ is the identity of the monoid operation. Consider the two-element poset $X=\left\{x,y\right\}$ with order $\leqslant\,=\mathrm{id}_{X}$. We let $E=X^{2}$ and define the order automorphism $\alpha=\left\{\left(x,y\right),\left(y,x\right)\right\}$. We get that $\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ has as its underlying lattice the 16-element Boolean lattice. It is straightforward to calculate that ${\leqslant}={\sim}{\leqslant}={\leqslant}^{c\smile}\circ\,\alpha$ and then to show that the set of binary relations $\{\varnothing,\leqslant,X^{2}\}$ is closed under $\cap$, $\cup$, $\circ$, $\Rightarrow$ and $\sim$}. Define a map $h:S_{3}\to\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ by $h(a_{-1})=\varnothing$, $h(a_{0})={\leqslant}$ and $h(a_{1})=X^{2}$. It is easy to check that this is indeed an embedding of $\mathbf{S}_{3}$ into the direct reduct of $\mathfrak{D}(\mathbf{E})$.
@@
$\left(p,q\right)\in\mathrm{id}_{\mathbb{Q}^{n}}\qquad$ $\textnormal{ iff }\qquad p=q$
$(p,q)\in L_{1}\qquad$ $\textnormal{ iff }\qquad p_{1}<q_{1}.$
@@
These relations are upsets of $\langle E,\preccurlyeq\rangle$ and form a Sugihara chain of length $2n+2$, i.e.,
@@ 3 Constructing finite odd Sugihara chains
The first lemma in this section establishes that when we have a poset $\langle X,\leqslant_{X}\rangle$ constructed from a chain as above, the order $\leqslant_{X}$ will be fixed by the $\sim$} operation described in Theorem 2.7. Note that $(x^{b},y^{d})\in{\sim}R$ if and only if $(y^{-d},x^{b})\notin R$. In particular, $(x^{b},y^{d})\in 0$ if and only if $y^{-d}\nleqslant x^{b}$.
@@
To prove the left-to-right inclusion, assume $x^{b}\leqslant_{X}y^{d}$. Then $x^{b}=y^{d}$ or $x<y$. If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{d}\right)\in 0$. On the other hand, if $x<y$, then $y\not\leqslant x$, so $y^{-d}\not\leqslant x^{b}$. Again we have $(x^{b},y^{d})\in 0$. For the other inclusion, let $\left(x^{b},y^{d}\right)\in 0=\left(\leqslant_{X}\right)^{c\smile}\circ\alpha$. Then we have $y^{-d}\not\leqslant x^{b}$. If b=db=-d, then $y\not\leqslant x$, so $x<y$. Hence, $x^{b}\leqslant_{X}y^{d}$. On the other hand, if $b\neq-d$, then $y^{-d}\not\leqslant x^{b}$ implies $y\not<x$. This means $x\leqslant y$, and so, since $d=b$, we get $x^{b}\leqslant_{X}y^{d}$.
@@
$\left(p,q\right)\in\mathrm{id}_{\mathbb{Q}^{n}}$  $\quad\textnormal{ iff }\quad p=q$
$\left(p,q\right)\in L_{1}$  $\quad\textnormal{ iff }\quad p_{1}<q_{1}$
@@
$\left(p,q\right)\in L_{i}$  $\quad\textnormal{ iff }\quad\left(p_{1},\ldots,p_{i-1}\right)=\left(q_{1},\ldots,q_{i-1}\right)\textnormal{ and }p_{i}<q_{i}.$
@@
Both $X^{2}$ and $\varnothing$ are upsets of $\left\langle X^{2},\preccurlyeq\right\rangle$. The fact that $\leqslant_{X}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ follows from Lemma 2.4. It would also follow from Lemmas 2.3 and 2.6 that $R_{i}$ is an upset for $i\in\left\{1,\ldots,n\right\}$, if we can show that $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ when $n\leqslant i<0$.
@@
Notice that when defining $R_{i}$ for $i\in\left\{1,\ldots,n+1\right\}$ we chose to use the unary order-reversing operation $\sim$ instead of -. Since $\alpha\neq\text{id}_{X}$ we know from Theorem 2.7 that the algebra will not be cyclic and hence ${\sim}R\neq-R$ in general. However, the following lemma shows that for the relations $R_{i}$, we will in fact get ${\sim}R_{i}=-R_{i}$.
@@
For each $i\in\left\{-n-1,\ldots,n+1\right\}$, we have $R_{i}={\sim}R_{i}$.
@@
Case 1: $n-1\leqslant i<0$. Let $\left(p^{b},q^{d}\right)\notin-R_{i}$. By Lemma 2.2, $\left(p^{b},q^{d}\right)\in\left(-R_{i}\right)^{c}=\left(\alpha\circ\left(R_{i}\right)^{c\smile}\right)^{c}=\alpha\;\circ\;R^{\smile}_{i}$. Therefore, $\left(q^{d},p^{-b}\right)\in R_{i}$, and so $\left(q^{d},p^{-b}\right)\in U_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. This implies $(q,p)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. It thus follows that $\left(q^{-d},p^{b}\right)\in U_{k}\subseteq R_{i}$, i.e., $\left(\alpha\left(q^{d}\right),p^{b}\right)\in U_{k}\subseteq R_{i}$. Hence, $\left(p^{b},\alpha\left(q^{d}\right)\right)\in R_{i}^{\smile}$, and therefore $\left(p^{b},\alpha\left(q^{d}\right)\right)\in\left(\left(R_{i}\right)^{c\smile}\right)^{c}$. From this we get $\left(p^{b},q^{d}\right)\in\left(\left(R_{i}\right)^{c\smile}\right)^{c}\circ\alpha=\left(\left(R_{i}\right)^{c\smile}\circ\alpha\right)^{c}=\left({\sim}R_{i}\right)^{c}$, i.e., $\left(p^{b},q^{d}\right)\notin{\sim}R_{i}$.
@@
Case 2: $0<i\leqslant n+1$. This case follows from the previous case and the fact that - and $\sim$ satisfy the involutive property.
@@
This is clearly true for $i=0$ and $i=n+1$. Since $\varnothing$ is vacuously transitive, the statement also holds for i=n1i=-n-1. We now consider the following two cases:
Case 1: $n\leqslant i<0$. Let $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. Then there exist $j,k\in\left\{1,\ldots,n+1+i\right\}$ such that $(p,r)\in L_{j}$ and $(r,q)\in L_{k}$. Hence, by Lemma 3.3, we have $\left(p,q\right)\in L_{\text{min}\{j,k\}}$, and so $\left(p^{b},q^{d}\right)\in U_{\text{min}\{j,k\}}\subseteq R_{i}$.
@@
If i=n1i=-n-1 (i.e. $R_{i}=\varnothing$) then we clearly have $R_{i}\circ R_{j}=R_{i}$. The case for $j=0$ follows from Lemma 2.2. So we consider the following cases:
Case 1: n1<i<j<0-n-1<i<j<0. Since $R_{j}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{i}$. Let $\left(p^{b},q^{d}\right)\in R_{i}$. Hence, there exists $k\in\left\{1,\ldots,n+1+i\right\}$ such that $\left(p^{b},q^{d}\right)\in U_{k}$, i.e., there exists $k\in\left\{1,\ldots,n+1+i\right\}$ such that $\left(p,q\right)\in L_{k}$. There are two cases:
@@
Case 2: n1<i<0-n-1<i<0, $j>0$. Since ${\leqslant_{X}}\subseteq R_{j}$, we get $R_{i}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. From the first part we get $\left(p^{b},r^{e}\right)\in U_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$, which means $(p,r)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. The second part gives $\left(r^{e},q^{d}\right)\in{\sim}R_{-j}=\left(R_{-j}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},r^{b}\right)\notin R_{-j}$. This implies that $\left(q^{-d},r^{e}\right)\notin U_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-j\right\}$. That is, $q_{1}\not<r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $q_{\ell}\not<r_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we get $r_{1}\leqslant q_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $r_{\ell}\leqslant q_{1}$. If $p_{1}<r_{1}$, then $p_{1}<r_{1}\leqslant q_{1}$, and so $\left(p,q\right)\in L_{1}$, i.e., $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}$. If $r_{1}<q_{1}$, we have $p_{1}\leqslant r_{1}<q_{1}$, so again $\left(p,q\right)\in L_{1}$, i.e., $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}$. If $p_{1}=r_{1}=q_{1}$, then it must be the case that $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}$.
Case 3: $i>0$, n1<j<0-n-1<j<0. Since $R_{j}\subseteq{\leqslant_{X}}$ we get $R_{i}\circ R_{j}\subseteq R_{i}$. Let $\left(p^{b},q^{d}\right)\in R_{i}$. Then $\left(p^{b},q^{d}\right)\in{\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},p^{b}\right)\notin R_{-i}$. This means that $\left(q^{-d},p^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. That is, $q\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{k-1}\right)=\left(q_{1},\ldots,q_{k-1}\right)$ then $q_{k}\not<p_{k}$. We therefore obtain $p_{1}\leqslant q_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{k-1}\right)=\left(q_{1},\ldots,q_{k-1}\right)$ then $p_{k}\leqslant q_{k}$. Suppose $p_{1}<q_{1}$. By the density of $\mathbb{Q}$, there is some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Now let $r=\left(p_{1},0,\ldots,0\right)$. Then $(p,r)\in L_{1}$ and $(r,q)\in L_{1}$, i.e., $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{-i}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{-i}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. On the other hand, suppose $p_{1}=q_{1}$. We have two cases: